Find the best investment and financing program using the Dean Model.

Dean Model Exercise Initial investment Project Option Amount Interest rate 400 19.1% 200 11% 2 200 18.5% 300 10% 3 100 300 12% 4 100 6.5% 200 15% 5 50 13% Find the best investment and financing program using the Dean Model
Answer #1

Dean Model is simultaneous analysis of investment and financing options. Mathematically the model can be formulated using the variables and parameters below:

Objective function:

\sum Aj1* Xj +\sum Di1* Yi \Rightarrow Maximize called as compounded Value (CV)

with \sum Aj0* Xj +\sum Di0* Yi = 0

here Xj is extent of investment projects (here 1 to 5)

and Yj is the extent of financing options (here 1 to 4 for A to D)

and Aj1 and Di1 are the net cashflows associated with each project in time 1 and 0 for O subscripts

Let us consider the below provided data:

Invesment Options Initial invesment IRR Rank Capital Flow at year 1
1 400 19.10% 2 476.4
2 200 18.50% 3 237
3 100 22% 1 122
4 100 6.50% 5 106.5
5 50 13% 4 56.5
Financing options Amount Interest Rank Capital flow at year 1
A 200 11% 2 222
B 300 10% 1 330
C 300 12% 3 336
D 200 15% 4 230

The capital flow for first yearin invesments and financing is calculated using following function= initial investment/amount *(1+ IRR or interest) and rank is derived on the basisof IRR and interest. (lower interest better and Higher IRR is better)

Hence the objective function stands for year 1 as:

476.4*x1 + 237* x2 + 122* x3 + 106.5* x4 + 56.5* x5 - 222* y1 - 330* y2 - 336* y3 - 230* y4 >>>>> Maximize

with for year 0 (initial values)

-400*x1 -200* x2 -100* x3 - 100* x4 - 50* x5 + 200* y1 + 300* y2 + 300* y3 + 200* y4

1. Eliminating invesmentes and financing options

invesment options and financing options can be eliminated first on the basis of graph plotting on the accumulated manner and in ranking order.

invesments options financing option Initial invesments interest rate of financing IRR of projects
3 2 50 10% 22%
3 2 100 10% 22%
3,1 2 150 10% 19.10%
3,1 2 200 10% 19.10%
3,1 2 250 10% 19.10%
3,1 2 300 10% 19.10%
3,1 2,1 350 11% 19.10%
3,1 2,1 400 11% 19.10%
3,1 2,1 450 11% 19.10%
3,1 2,1 500 11% 19.10%
3,1,2 2,1,3 550 12% 18.50%
3,1,2 2,1,3 600 12% 18.50%
3,1,2 2,1,3 650 12% 18.50%
3,1,2 2,1,3 700 12% 18.50%
3,1,2,5 2,1,3 750 12% 13%
3,1,2,5,4 2,1,3 800 12% 6.50%
3,1,2,5,4 2,1,3,4 850 15% 6.50%
2,1,3,4 900 15%
2,1,3,4 950 15%
2,1,3,4 1000 15%

From the above accumulated numbers it is clear that invesment option 4 and financing option 4 (D) are completely out of picture as at this point the interest fall above IRR of projects.

hence options left for invesments are 1,2,3 and 5. and financing options left are 1,2,3.

The possible combinations of invesments with these options are as follows with their compounded values:

proportion of funding type Compounded value
Strategy Number invesments considered invesments required 1 2 3
1 1 400 50% 100% 0% 35.4
2 2 200 0% 67% 0% 17.22
3 3 100 0% 33% 0% 12.11
4 5 50 0% 17% 0% 1.5
5 1,2 600 100% 100% 33% 49.512
6 1,3 500 100% 100% 0% 46.4
7 1,5 450 75% 100% 0% 36.4
8 2,3 300 0% 100% 0% 29
9 2,5 250 0% 83% 0% 18.5
10 3,5 150 0% 50% 0% 13.5
11 1,2,3 700 100% 100% 67% 59.624
12 2,3,5 350 25% 100% 0% 30
13 3,5,1 550 100% 100% 17% 46.9
14 1,2,5 650 100% 100% 50% 49.9
15 1,2,3,5 750 100% 100% 83% 59.9

Hence from the above options , strategy number 15 is the best option as it has highest Coompounded Value of 59.9

hence best option with dean model is investment in 1,2,3,5 with funding from A, B and C (83.33%).